A horizontal force of 150 N is used to push a 38.0 kg packing crate a distance of 6.85 m on a rough horizontal surface. If the crate moves at constant speed, find each of the following.

a) find the work done by the 150N force:

b) find the coefficient of kinetic energy between the crate and rough surface
Answer #1

Force F = 150N, distance s = 6.85m, mass m=38.0kg

a)

the work done by the 150N force is:

W1 = F.s = 150N*6.85m = 1027.5J

b)
let the the coefficient of kinetic energy between the crate and rough surface be
μk, since it move in constant speed, acceleration a = 0

F - mgμk = ma = 0

μk = F/mg = 150N / (38.0*9.8m/s2) = 0.40

the the coefficient of kinetic energy between the crate and rough surface is 0.40